OpenStudy (anonymous):

what two numbers, when multiplied, give a product of -160; yet when added, equal 6?

OpenStudy (jamesj):

So let the two numbers be x and y. Then xy = -160 and x + y = 6 From the second equation, we have that y = 6 - x Substitute that now into the first equation and solve for x.

OpenStudy (anonymous):

xy = -160 x + y = 6 Now evaluate y from one and put it into other to get x value..

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

Try with factors of -160.. 8*(-20) --------- sum is not 6 10*(-16) -----------sum is not 6 16*(-10) Yes this produces -160 and sum = 16 - 10 = 6

OpenStudy (anonymous):

Yes, thanks JamesJ (& waterineyes).